OpenStudy (anonymous):

Medal+fan! Help solving pleeaseeee. (-2 - x)^0.5 - x = 2

hartnn (hartnn):

(-2 - x)^0.5 - x = 2 (-2 - x)^0.5 =x +2 now square both sides, what u get ?

OpenStudy (tanner23456):

-2-x = (x+2)^2 -2-x = x^2+4x+4 combine like terms x^2+5x+6=0 now factor (x+2)(x+3) now set each equal to zero x+2 = 0 x+3 = 0 x= -2,-3

hartnn (hartnn):

'x' can't be -3

OpenStudy (tanner23456):

True, good catch.

OpenStudy (anonymous):

make sure whenever you solve radical equations you check for extraneous solutions...